OpenStudy (konradzuse):

integral of 3x^2 * sin(x^3)dx

OpenStudy (konradzuse):

I did\[ u = x^3\] \[du = 3x^2 dx\] \[\int\limits \sin(u) du\] -cos(u) + c \[-\cos(3x^2) + c\] but apparently that is wrong?

OpenStudy (konradzuse):

Oh wait... it should be -cos(x^3) + c

myininaya (myininaya):

yes because u equals x^3 not 3x^2 :)

OpenStudy (konradzuse):

Yeah stupid mistake late at night :D.

myininaya (myininaya):

it happens to all man